### v1

# Mobile Edge Generation-Enabled Digital Twin: Architecture Design and Research Opportunities

Abstract
A novel paradigm of mobile edge generation (MEG)-enabled digital twin (DT) is proposed, which enables distributed on-device generation at mobile edge networks for real-time DT applications. First, an MEG-DT architecture is put forward to decentralize generative artificial intelligence (GAI) models onto edge servers (ESs) and user equipments (UEs), which has the advantages of low latency, privacy preservation, and individual-level customization. Then, various single-user and multi-user generation mechanisms are conceived for MEG-DT, which strike trade-offs between generation latency, hardware costs, and device coordination. Furthermore, to perform efficient distributed generation, two operating protocols are explored for transmitting interpretable and latent features between ESs and UEs, namely sketch-based generation and seed-based generation, respectively. Based on the proposed protocols, the convergence between MEG and DT are highlighted. Considering the seed-based image generation scenario, numerical case studies are provided to reveal the superiority of MEG-DT over centralized generation. Finally, promising applications and research opportunities are identified.
Index Terms:
Artificial intelligence generated content (AIGC), digital twin, mobile edge computing, mobile edge generation.
I Introduction
With recent decade’s booming developments in artificial intelligence (AI) and wireless networks, digital twin (DT) has emerged as a promising approach for creating virtual replications of real-world physical objects and processes [1]. By revolutionizing the fashions of modelling, monitoring, analytics, and decision-making for intricate real-world systems, DT is anticipated to transform diverse facets of human society, spanning industry, education, agriculture, healthcare, and beyond. To maintain immersive and real-time evolving virtual duplicates for physical entities, DT requires ubiquitous connectivity, high-speed transmission, low latency, and intensive computing on mobile devices [2, 3]. To meet these challenges, a prospective solution is to integrate DT with the forthcoming sixth generation (6G) wireless networks, which leads to a win-win option. On the one hand, 6G wireless networks enable real-time synchronized DTs by incorporating new communication technologies (e.g., millimetre-wave/terahertz communications and satellite communications), enhanced infrastructures (e.g., integrated sensing and communications), as well as cloud/edge intelligence [3]. On the other hand, DT provides powerful tools to realize precise modelling (e.g., channel modelling) and self-sustaining management of 6G wireless networks [4].
While holding immense potentials, establishing and maintaining high-fidelity virtual twins suffer from serious challenges and significant costs, particularly for real-time mobile DT applications. Fortunately, benefiting from advances in generative AI (GAI) and large-scale Transformer models, AI generated contents (AIGC) has unlocked new pathways to generate high-fidelity and vivid digital contents in virtual worlds [5]. The great success of GAI models, such as Generative Pre-training (GPT) [6] and Stable Diffusion [7], has spawned numerous AIGC applications in Natural Language Processing (NLP) and computer vision, involving text generation, text-to-image generation, image translation/synthesis, and video generation. With the impressive capabilities to comprehend data patterns and mimic human creations, AIGC can automatically and seamlessly map real-world contents into virtual twins. Therefore, AIGC has been regarded as a key enabler of DT for smart content transformation and enhancement [5]:
  • Content Transformation: Constructing virtual twins of physical worlds requires seamless transformations across multi-modal data (e.g., multilingual text, audio, image, video, and 3d point clouds) via AIGC.
  • Content Enhancement: The completeness, quality, and customization of digital contents should be improved by AIGC. For instance, AIGC can compensate for missing information, augment image/video quality, mitigate distortion/noise, edit content via human-computer interaction, and so on.
Despite these promising benefits, the size of GAI models has increased to 175 billions of parameters in GPT-3 from several hundreds of millions of parameters in GPT [6]. Following the stunning generative performance presented by GPT-3-style models, the development of large models has become a popular pursuit in AIGC domain. In the most recent GPT-4 model, the number of parameters has grown to an unprecedented 1.7 trillion, which is nearly 1000 times larger than GPT-3. However, training and executing a large-scale GAI model rely on high computing powers from massive GPUs [8]. Therefore, the deployments of large-scale GAI models on mobile devices suffer from overwhelming hardware complexity and serious energy concerns, which limits their applicabilities in real-time DT. To relieve UEs’ computational burdens, mobile edge networks can be deployed for DT, where the generation tasks can be offloaded from hardware-constrained user equipments (UEs) to high-performance edge servers (ESs). However, as massive mobile devices may require AIGC capabilities in real-time DT applications, transmitting huge-volume data for centralized generation is time-consuming and privacy-sensitive. This requires edge learning and inference design for large-scale GAI models in DT systems.
Recently, the concept of mobile edge generation (MEG) has been proposed [9]. Specifically, MEG decentralizes GAI models into sub-models deployed on UEs and ESs. By doing so, distributed generation can be achieved by transmitting intermediate features of GAI sub-models between the heterogeneous hardware devices, without exchanging the raw data. In comparison to centralized generation, MEG can significantly reduce the latency and hardware costs for mobile UEs to access large-scale GAI models. Therefore, it enables an innovative paradigm to exploit large-scale GAI models like GPT-4 and DALL-E in real-time DT systems. However, to the best of our knowledge, the integration of MEG and DT has not been explored yet. To fill in this gap, we propose a novel MEG-enabled DT (MEG-DT) architecture in this article, targeted at realizing efficient on-device generation at the mobile edge networks of DT. Our main contributions can be summarized as follows.
  • We propose a novel MEG-DT architecture, which decentralizes large-scale GAI models onto heterogeneous hardware devices at mobile edge networks. The proposed MEG-DT can provide low-latency, privacy-preserving, and individual-level generation for real-time DT services, thus addressing the limitations of centralized generalization. We discuss the key advantages and conceive single-user and multi-user generation mechanisms.
  • We explore two operating protocols in MEG, namely sketch-based generation and seed-based generation. To achieve efficient distributed generation, these protocols transmit interpretable and latent features between heterogeneous hardware devices, respectively. We highlight that the convergence of MEG and DT can be realized, which empowers autonomous DT operation and self-optimized MEG configuration.
  • We identify suitable generation mechanisms and protocols for image, video, and interactive-scene generations in MEG-DT. We further provide numerical case studies for seed-based image generation in MEG-DT, which demonstrates that the proposed architecture achieves high-quality and noise-resistant generation compared to centralized generation even in low signal-to-noise (SNR) conditions, whilst significantly reducing overheads.
II MEG-DT: Architecture, Key Features, and Mechanisms
Figure 1: The vision of the proposed MEG-DT architecture.
In this section, we first propose a novel MEG-DT architecture. Then, we highlight the key advantages of MEG-DT and present tailored single-user and multi-user generation mechanisms.
II-A The Proposed MEG-DT Architecture
As shown in Fig. 1, we consider a mobile edge network for MEG-DT that consists of heterogeneous hardware devices, including high-performance ESs and low-cost UEs. To realize on-device mobile generation, large-scale GAI models are decentralized into collaborative GAI sub-models that are jointly deployed on ESs and UEs. By transmitting the features extracted from GAI sub-models between heterogeneous hardware devices, high-quality generation can be realized by jointly exploiting datasets, computational resources, and customized sub-models on both sides. In this way, MEG can bridge between ESs and mobile UEs as an immediate layer for DT construction and applications. With the tailored deployment and design of the distributed GAI sub-models, the proposed MEG-DT architecture can realize real-time content translation and enhancement to provide on-device DT services, such as multi-modal sensing, virtual replica construction, data analytics, performance prediction, and decision-making.
II-B Key Advantages of MEG-DT
The key advantages of MEG can be summarized as follows.
  • Low-Latancy Feature Compression and Transmission: Unlike centralized generation that directly exchanges raw data between UEs and ESs, MEG transmits features constructed from sub-models, which encode and compress the raw data into low-dimensional tensors. Therefore, the communication overheads and end-to-end generation latency required by MEG can be significantly reduced.
  • Distributed Model Deployment and Generation: MEG enables the distributed deployment of a large-scale GAI model on mobile UEs and the high-performance ES. With this innovative paradigm, the computing resources and local datasets from heterogeneous hardware devices can be efficiently orchestrated in various application scenarios.
  • Privacy Preservation and Security: Since UEs/ES can selectively expose and share their private data with each other, MEG can preserve the data privacy for various DT applications. Furthermore, by transmitting the implicit features during content generation, the system security can be significantly improved.
  • Individual-Level Customized Generation: Relying on the MEG paradigm, the model parameters, hyperparameters, and neural layer structures of GAI sub-models deployed on different UEs can be customized according to their individual preferences, particular generation purposes, and specific hardware constraints.
II-C Single-User and Multi-User Generation for MEG-DT
Based on the proposed MEG-DT architecture, we discuss the single-user and multi-user generation mechanisms.
Figure 2: Single-user generation mechanisms for MEG-DT.
II-C1 Single-User Generation
As shown in Fig. 2, the proposed MEG-DT architecture mainly contains four single-user generation mechanisms. Specifically, UE-to-ES (U2E) generation and ES-to-UE (E2U) generation enable feature encoding and content generation at ES and UE, respectively. Furthermore, sequential ES-and-UE (SEU) generation and parallel ES-and-UE (PEU) generation enable feature encoding and content generation on both sides, as discussed as follows.
  • U2E Generation: By sensing physical environment, UE encodes the collected data via its GAI sub-model to construct low-dimensional features. After receiving features transmitted from UE, ES generates digital contents via its GAI sub-model. A typical use case of U2E generation in DT is to generate/refresh the virtual replica at ES using the multi-modal sensing data (e.g., text, voice, camera, radar/LiDAR, inertial measurement unit (IMU)) from UE.
  • E2U Generation: ES encodes the local data and/or the inference results via its GAI sub-model and transmits the encoded features to target UE. By decoding the received features, UE further reconstructs inference results and generates desired contents. For example, UE can generate desired predictions by retrieving data (e.g., high-precision 3D map and expert knowledge) from the server.
  • SEU Generation: SEU generation allows ES and UE to sequentially transmit encoded features and generate contents, enabling virtual-reality synchronization. UE first transmits extracted features to ES in the uplink. Next, ES generates contents and updates server dataset, and then transmits downlink features to assist in the generation on UE. This mechanism can be employed when UE has very limited computing/storage capacities and requires ES to handle the vast majority of computing tasks.
  • PEU Generation: PEU generation allows UE and ES to simultaneously perform feature encoding, transmission, and content generation in a parallel manner, as shown in Fig. 2(d). Specifically, a feature gate is learned to integrate features from both sides. Local copies of the feature gate are deployed on both ES and UE, which create representations of a synchronized and shared view for parallel generation. Compared to SEU generation, this mechanism can reduce end-to-end generation latency for delay-sensitive DT applications, but requires higher hardware costs for UEs.
II-C2 Multi-User Generation
Since each UE can only perceive their surrounding environment, the application of MEG-DT may require interactions between multiple UEs. By integrating UEs’ local features, a global view of physical/digital world can be established for immersive generation and cooperative decision-making. To reap the benefits, we conceive three multi-user generation mechanisms for MEG-DT, which can realize distinct levels of interactions and coordination between UEs, as shown in Fig. 3. Here, we focus on multi-user interactions for UE-side generation. The BS-side interaction can be achieved similarly.
Figure 3: Multi-user generation mechanisms for MEG-DT.
  • Individual Generation via Local Features: In the first level, UEs independently perform single-user generation using their local features. As shown in Fig. 3(a), each UE utilizes its sensors to collect local environmental information and individually interacts with the ES. Thus, a real-time and local view of physical world can be acquired and projected onto the 3D map. This empowers fast generation and decision-making for low-cost UEs.
  • Decentralized Generation via Shared Features: The individual generation performance is limited by the local view of the real-world environment observed by each UE. As a remedy, the second level envisions UEs to dynamically share their encoded features. By integrating features shared by nearby UEs, each UE can asynchronously perform decentralized generation once the local observation is renewed, thus enabling multi-device interactions with a low latency.
  • Coordinated Generation via Fused Features: The third level enables UEs to thoroughly fuse their local features at the nearby ES for enhanced coordination. Exploiting the fused features, fully synchronized and coordinated multi-user generations can be realized. This mechanism can be employed when UEs require a comprehensive picture of physical world, while tolerating increased latency and computing complexity.
III MEG-DT: Principle and Interplays
To realize efficient distributed generation during the single-user and multi-user generations, a key issue is to effectively extract low-dimensional features that encode local data at specific ESs/UEs. In this section, we will introduce the main principles to extract useful features by exploring two protocols, i.e., sketch-based generation and seed-based generation. Then, the convergence between MEG and DT will be discussed.
III-A Protocol of MEG: Sketch-based or Seed-based?
Depending on whether the transmitted features can be explicitly interpreted or not, we introduce two basic operating protocols for MEG, namely sketch-based generation and seed-based generation, as discussed as follows.
III-A1 Sketch-Based Generation
Sketch typically represents a simplified image or drawing that only retains object outlines and essential spatial characteristics of the original image. As it provides a low-complexity visual representation of image, it has been widely applied in image retrievals and analytics. Derived from this concept in computer vision, in the considered MEG-DT settings, we refer to the sketch as a low-complexity and interpretable abstraction of the raw data (such as text, image, and 3D maps), which can be used for a wide rage of generation tasks.
To efficiently perform sketch-based generation, ESs and UEs can jointly train domain-specific encoder and decoder that bridge between the distributed GAI sub-models, as shown in Fig. 4(a). Specifically, the encoder learns to extract a sketch of the raw data, thus preserving important information while reducing data volumes. Then, by transmitting the obtained sketch to the decoder deployed on another hardware equipment, the details will be reconstructed for high-quality generations. To efficiently compress and reconstruct data, the encoder and decoder can inherit the architectures from current GAI models for image-to-image translation [10], e.g., generative adversarial networks (GANs), variational auto-encoder (VAE), and diffusion models. These models can be further tailored for different generation purposes. During generation, text prompts can be combined to control the content editing and synthesis, thus flexibly synthesizing virtual/realistic datasets on ES and UEs. In general, sketch-based generation is easy to implement, guide, and examine.
Figure 4: Diagram of sketch-based generation and seed-based generation protocols.
III-A2 Seed-Based Generation
Instead of transmitting sketches that have explicit interpretations, another promising approach is to transform the high-dimensional data into latent features without intuitive explanations. These latent features are known as seeds. In comparison to the sketch-based generation that creates contents on the original feature space of the physical-world data, seed-based generation can deal with a more compact latent feature space. By identifying the underlying dependence of data implicitly, seed-based generation is more efficient to process large-scale or sequential data, such as time-varying sensing observations and videos.
A simple implementation of seed-based generation is to split an existing GAI model into two sub-model. Thus, seeds can be constructed by the hidden layer that splits the original GAI model. During forward/backward propagation, only features/gradients of the split hidden layer need to be transmitted between different devices. In practice, pretrained GAI models (e.g., Stable Diffusion) can be directly split into encoder and decoder for seed-based generation without additional training process. Nevertheless, the achievable end-to-end generation quality is still limited by practical compression performance and noisy wireless channels. To fully exploit the potentials of seed-based generation and overcome these limitations, advanced MEG-DT strategies should be developed to systematically realize efficient seed construction, compression, and source-channel coding, thus enhancing on-device generation quality with lower system complexity and overheads, as discussed as follows.
  • End-to-End Coding and Transmission: To enhance the end-to-end MEG performance, GAI sub-models can be designed to accomplish joint source-channel coding (JSCC) for transmitting seeds in noisy wireless channels. As shown in Fig. 4(b), at seed transmitters, GAI sub-models learn to directly map the source signals of the constructed seeds into channel symbols. Then, at seed receivers, seeds are further estimated from noisy channels to generate desired contents. The distributed GAI sub-models can be jointly trained to enhance the resulting content generation quality, whilst minimizing seed distortions. This can efficiently mitigate channel noises and potential hardware impairments for seed transmissions.
  • Seed and Sub-Model Compressions: To enable lightweight computing and low-latency generation, it is crucial to learn tiny GAI sub-models and acquire dimension-reduced seeds in MEG-DT. To fully reduce generation overheads exploiting latent features, both seeds and GAI sub-models can be efficiently compressed. Specifically, the dimensions of seeds constructed by distributed sub-models should be adaptively lessened by removing the insignificant neurons and information via pruning. Furthermore, the GAI sub-models should be flexibly slimmed into tiny neural structures based on dynamic distillation [11] or truncated diffusion, thus reducing computational burdens on UEs.
Figure 5: The suitable MEG protocols for different generation tasks.
III-B Convergence between MEG and DT
The proposed MEG-DT architecture empowers the convergence and interplays between MEG and DT, which can be analysed as follows.
  • MEG Aided DT: MEG enables autonomous creation, maintenance, and validation of DT on mobile devices. Specifically, the decentralized GAI sub-models can be trained to minimize the divergence between the digital model distribution and the real target distribution, thus accurately modelling the real-world complex systems to construct high-fidelity digital replications. Notably, MEG can creatively generate vivid images, videos, and voices in the virtual world by synthesizing the inputs from different mobile UEs. Furthermore, MEG can learn to emulate the expert behaviours in the high-fidelity virtual world for decision-making. Additionally, testing scenarios can be generated via MEG for validation, thus assisting the online self-prediction, self-diagnose, and self-management of DT systems.
  • DT Aided MEG: A virtual twin can be created by DT to capture the dynamics of mobile edge networks based on data analytics and multi-physics simulation. This can enable self-optimized configuration of MEG. Specifically, DT can simulate the real-world channel models and track UEs’ mobility, thus empowering low-overhead and high-reliability sketch/seed transmissions for MEG. Furthermore, based on the predictive analytics of UEs’ behaviours and environmental changes, reliable and secure distributed computing schemes for MEG can be derived in DT, which balances the computation/communication workloads and diagnoses potential malicious attacks of mobile devices.
IV Case Studies
In this section, we identify suitable generation mechanisms/protocols for multi-modal DT generation and demonstrate the performance of E2U seed-based image generation via numerical examples.
IV-A Multi-Modal Generation
DT involves generations of multi-modal contents, including text, images, videos, and interactive scenes, etc. We summarize suitable protocols for different generation tasks in Fig. 5. Text generation can be performed via centralized GAI models as it requires only a small transmission bandwidth. Here, we discuss image/video generation and interactive scene generation.
  • Image Generation: Conventional image generation tasks often perform generation based on prompts. The prompts can be text, image, or even multi-modal data. Regarding text-to-image generation, the bandwidth required for text transmission is much less than that of image. Hence, to reduce the communication overheads, it is suitable to employ the E2U/U2E mechanism for text-to-image generation at UEs/ESs. Regarding image-to-image generation, SEU/PEU mechanism can be employed for sequential or synergetic feature encoding. On the other hand, both seeds and sketches are feasible for conveying compressed information, but have distinct advantages. The seed can achieve higher compression rates, while sketch can provide human-readable visual previews.
  • Video Generation: The video transmission requires enormous bandwidth, especially when high-resolution virtual worlds should be constructed. For image-to-video generation, applying SEU protocol for video generation can efficiently reduce the size of the video file. Nevertheless, the users can only receive the generated video after the generation process, which may increase the waiting time for users. In this case, the PEU generation has a chance to provide a simultaneous generation and presentation of the video. To enhance compression ratio, seed-based generation is preferable in this scenario.
  • Interactive Scene: A key function of DT is to provide interaction. Thus, the generation of interactive scenes is a crucial part of MEG-DT. Compared with image/video generation, interactive scenes have higher requirements for latency. Therefore, the PEU protocol should be adopted in the generation of interactive scenes for virtual-reality convergence. In addition, as interactive scenes require continuous visualization results (even with insufficient clarity), sketch is more suitable for serving the generation of interactive scenes.
(a) Generation quality.
(b) FID score.
Figure 6: Performance comparisons for generation approaches.
IV-B Performance Evaluation
In this part, we provide case studies of text-to-image generation for the proposed MEG-DT architecture. The seed-based E2U generation protocol is investigated under the compact additive white Gaussian noise channel. To train distributed GAI sub-models for seed encoding/decoding, images generated by centralized GAI and transmitted under perfect channel conditions are set as the ground truth. Given the ground truth, we calculate the Fréchet Inception Distance (FID) score to measure the distortion of reconstructed images via MEG-DT. We assume $3232$-bit floating point number for transmissions. The following three generation approaches are considered in the simulation:
  • Centralized Generation: A pretrained Stable Diffusion model [7] is utilized as baseline for centralized generation. An image with $33.5533.55$ Mbits is generated by ES and is transmitted to UE in the downlink.
  • MEG-DT: By splitting the pretrained Stable Diffusion model, a denoising diffusion implicit models (DDIM) sampler is deployed on ES, which samples a $0.520.52$-bit latent feature as a seed for distributed generation on UE.
  • E2E-MEG-DT: ES and UE train GAI sub-models in noisy wireless channels to achieve end-to-end coding and transmission of $0.260.26$-bit seeds. A JSCC encoder/decoder is integrated into Stable Diffusion model for mapping seeds’ source signals into channel symbols.
Fig. 6 presents the numerical results, where a $1024\times 10241024 × 1024$ flower scene is generated by text prompt in the virtual twin. By transmitting seeds instead of images through the noisy wireless channels, MEG-DT schemes can achieve noise-resistant generation compared to centralized generation, whilst significantly reducing overheads. Furthermore, the E2E-MEG-DT scheme is more robust and exhibits a significantly higher gains in both generation quality and FID score, even in low SNR conditions. It is also worth noting that under high SNR conditions (SNR>20absent20>20> 20), additional slight distortion is introduced due to the difficulty in perfectly training the end-to-end encoder and decoder.
V MEG-DT: Applications and Open Problems
V-A Promising Applications of MEG-DT
We introduce several important applications of MEG-DT, which can be exemplified as follows.
V-A1 Smart City
: Smart city involves densely deployed sensors, interconnected vehicles/robots, social networks, and cellular/WiFi networks. With the huge amounts of data continuously produced by these entities, virtual city twins can be constructed via MEG-DT to integrate observations of massive mobile devices. This will empower accurate time-series traffic forecasting, social event forecasting, and rational decision-making for city governors, traffic administrator, and network operators, etc.
V-A2 Industry 4.0
: Industry 4.0 requires cyber-physical integration for smart manufacturing [13]. MEG can easily build cyber models of products using prompts and sensing data, thus achieving personalized and low-cost product designs. During manufacturing process, on-device virtual twins of smart factories can be created, thus supporting 24/7 real-time monitoring and quality control.
V-A3 Extended Reality (XR) and Metaverse
: The distributed GAI sub-models can synthesize diverse datasets collected by different devices, thus realizing the convergence between virtual and real worlds to build XR and metaverse applications. For instance, the vehicular XR metaverse can be built by synthesizing massive traffic and driving data, thus improving robustness, driving safety, traffic control efficiency, and sustainability [14].
V-A4 Healthcare
: Enabling mobile AIGC in DT is promising to provide personalized healthcare [15]. MEG-DT delivers a low-latency distributed solution for mobile AIGC, which can simulate data of rare diseases, establish testbeds for comprehensive drug validations and analytics, and provide ubiquitous personalized medical services.
V-B Open Problems of MEG-DT
While MEG-DT providing a promising distributed generation paradigm, there are still some open research problems in practical designs, which can be outlined as follows.
  • Security: Although seed-based generation enables privacy preservation, security remains an open challenge in MEG-DT especially for sensitive applications. Specifically, when the sketches/seeds are attacked or corrupted by poisoning, the content generation process may be collapsed. Furthermore, since sketches and seeds generally contain important object features and substantial semantic information, some privacy-sensitive information (e.g., local sensing data) may be eavesdropped by attackers through reverse engineering techniques, leading to privacy disclosure. These vulnerabilities pose an urgent requirement to incorporate advanced security mechanisms into the MEG-DT architecture, such as differential privacy and blockchain techniques.
  • Scalable Learning and Compression: There are several candidate techniques for the compression of sketch/seed and GAI sub-models [12]. Specifically, to achieve a compact model representation, inessential components can be removed by pruning, network sparsification, and knowledge distillation algorithms, thus shrinking the large-scale models with achievable performance loss. Another approach is to decompose cumbersome learning parameters into small-sized tensors via tensor decomposition, which can relieve memory requirements and computational complexity. Instead of reducing the number of model parameters, the quantization method focuses on achieving low-bit features. Note that the compression and transmission of learning features on UEs are simultaneously subject to the hardware complexity limitation and the radio resource scarcity. This constitutes a research interest of developing scalable learning and compression techniques that can balance between compression ratio, hardware applicability, and transmission quality (e.g., sketch/seed distortion, latency and reliability).
  • Distributed Computing and Coordination: MEG-DT depends on efficient distributed computing and coordination. On the one hand, the distributed deployment of GAI sub-models requires adaptive configuration to achieve the optimal performance with acceptable latency and complexity. For individual generation, the combined use of sketches and seeds should be designed to achieve extreme generation quality. For coordinated generation, UEs should dynamically determine the optimal set of interacting devices and the optimal frequency for sketch/seed interactions. On the other hand, the sensing, communication and computing resources over multiple devices should be coordinated to accurately track environment changes to refresh digital replica, as well as automatically making decisions in response to both current conditions and predictive changes. This requires the coordination between heterogeneous hardware devices and functionalities.
  • Personalization and Customization: Personalization is a crucial aspect for MEG-DT to meet users’ individual expectations. For personalized generation, GAI sub-models need to be customized on individual UEs that experience different data distributions, performance metric requirements, and hardware/software limitations. The personalized customization can be mainly achieved via two approaches, i.e., data augmentation and model adaptation. On the one hand, data augmentation directly tackles the data distribution drift, which pretrains the globally shared GAI sub-models with heterogeneous data and then fine-tunes the pretrained sub-models via UE’s local data. On the other hand, model adaptation achieves enhanced and accelerated personalization of GAI sub-models via few-shot learning (e.g., meta-learning and transfer learning) or knowledge transfer techniques. Specifically, meta-learning learns an optimal GAI meta-model that can fast generalize across a range of tasks by maximizing the model parameter sensitivity. Based on the meta-model, personalized sub-models for particular device can be achieved via few-shot adaptation. Moreover, transfer learning focuses on diminishing the domain discrepancy between the globally shared sub-models and personalized sub-models. Instead of transferring model parameters, knowledge transfer further shares knowledge between initial sub-models and personalized sub-models via soft predictions, which enables heterogeneous sub-model structures on different UEs.
VI Conclusion
A novel MEG-DT architecture has been proposed in this article for on-device DT generation at mobile edge networks. The key advantages of MEG-DT has been discussed, and single-user and multi-user generation mechanisms have been put forward. We introduced sketch-based and seed-based generations and discussed the convergence between MEG and DT. Furthermore, numerical case studies for the seed-based text-to-image generation have exhibited the superiority of MEG compared to centralized generation. The potential applications and open research problems have been identified.